# envname-P envname#1

### v2

@@ 5.1.1. Data Description
We conduct experiments on an urban travel itinerary dataset from four Chinese cities, collected in collaboration with a leading travel agency specializing in single-day city tours. The dataset comprises top-rated 1233 urban itineraries and 7578 POIs. Each data sample contains a user request, the corresponding urban itinerary plan, and detailed POI data. For an in-depth understanding of the data format, sample entries, and key preprocessing methodologies employed, refer to Appendix B.
5.1.2. Implementation Details
We utilize two LLMs through their APIs: GPT-3.5 (gpt-3.5-turbo version) and GPT-4 (gpt-4-turbo version). GPT-4 is used for the final itinerary generation ($\small\mathbb{P}_{IG}$) to guarantee the overall quality of the itinerary, and GPT-3.5 is employed for other LLM interactions to ensure the response speed of our OUIP service. We also incorporate the text-embedding-ada-002 model for embedding purposes. The spatial coherence of itineraries is optimized through an open-source TSP solver [^1] https://github.com/fillipe-gsm/python-tsp. The integration of POI data encompasses various attributes, including geographical coordinates, user ratings, categorizations, and physical addresses, facilitated through the Amap API [^2] https://lbs.amap.com/. For the production environment, our infrastructure leverages Amazon Elastic Container Service (ECS) to orchestrate a cluster of servers, each provisioned with an 8-core CPU and 16GB of RAM, optimizing performance for Python-based applications and machine learning APIs. Notably, the original languages of our service and data are simplified Chinese, and the translated version is used for demonstration purposes in this paper. More implementation details can be found in Appendix C.
[…]
Ours w/ GPT-3.5 27.6 79.4 0.56 67.3 58.8 61.4
Ours (full) 31.4 86.0 0.42 69.8 64.6 72.0
Table 2. Ablation study on Shanghai dataset.
5.2. Evaluation Metrics
According to the concept of the OUIP task, a satisfactory itinerary must possess the following characteristics: spatial coherence and alignment with the user’s needs. To this end, we have designed several objective metrics to evaluate the generated itineraries.
5.2.1. Rule-based Metrics
@@
We first design some rule-based metrics: (1) Recall Rate (RR): the recall rate of POIs in the ground truth itinerary; (2) Average Margin (AM): The average difference per POI between the total distance of the generated itinerary and the shortest distance (calculated by TSP); (3) Overlaps (OL): the number of self-intersection points in the generated itinerary, e.g. there are 2 overlaps in Fig. 3 (right). (4) Fail Rate (FR): The percentage of POIs generated by LLM that cannot be matched with the queried POIs from the map service (serve as the full POI set). RR evaluates the method’s accuracy in understanding user requests and recommending correct POIs. AM and OL measure spatial optimization for POI visit order, where lower is better. FR evaluates LLM’s information accuracy and indicates the method’s compatibility with the online service. Locations of failed POIs are inaccessible and cannot be displayed on the map, which lowers the user experience.
5.2.2. GPT-based Metrics
@@
The above rule-based metrics are intuitive. Some other aspects are hard to quantify, e.g. the intrinsic appeal of the POIs or the degree of alignment between user requests and the generated itinerary. Therefore, following previous benchmarks for automatic evaluation (Li et al., 2023; Bang et al., 2023; Wang et al., 2023b), we propose several GPT-based metrics: (1) POI Quality (PQ): how interesting and diverse the POIs are; (2) Itinerary Quality (IQ): the overall quality and coherence of the itinerary; (3) Matchness (Match): the matchness between the itinerary and the user request; We ask GPT-4 to rank two candidate itineraries generated with different methods and compute the win rate. We repeat the same query for at least 10 times with different seeds to obtain a reliable result.
Automated evaluation through GPT offers a more efficient and cost-effective alternative to human evaluation for large-scale experiments. For example, the cost of GPT tokens for the entire ablation study in Sec. 5.4 is less than \$25. In addition, the results of the online subjective evaluation in real-world scenarios in Sec. 5.6.2 demonstrate consistent results between human and GPT assessments.
5.3. Overall Results
5.3.1. Baselines
We compare ItiNera with the following baselines:
@@ 5.3.1. Baselines
  • IP (Gunawan et al., 2014): A traditional IP method. We simplified it to use LLM for time budgeting and to consider POI ratings as utilities.
@@
  • Ernie-Bot 4.0 (Sun et al., 2021): The best-performing model on Chinese LLM tasks, selected as our dataset and service are in Chinese.
  • GPT-3.5 (OpenAI, [n. d.]): The standard ChatGPT model.
  • GPT-4 (OpenAI, 2023): The enhanced model offering superior performance and broader knowledge coverage.
  • GPT-4 CoT (Wei et al., 2022): The model improves GPT-4 with the LLM-as-agent method, Chain-of-Thought reasoning.
We employ the same itinearry generation prompt for all baselines, including basic task requirements and output format as in $\small\mathbb{P}_{IG}$. Notably, for GPT-4 CoT, we extend the prompt by integrating the “thoughts”. These prompts are detailed in Appendix F.5. More details on baseline implementations can be found in Appendix C. Note that the baseline IP and our method does not compute Fail Rate as the candidate POIs are from the dataset.
5.3.2. Results Analysis
@@
We provide analysis for results (RQ1). As illustrated in Tab. 1, our proposed ItiNera significantly outperforms all baselines across all metrics. For rule-based metrics, ItiNera shows ~30% improvement over the best baseline, highlighting its superior capability in personalizing user experiences. Furthermore, it maintains spatial coherence by generating itineraries that are only marginally longer (~100 meters per planned POI) than the shortest path solved by TSP, a remarkable achievement considering potential user requests that may deviate from spatially optimal routes. Notably, ItiNera is the only method to surpass the performance of GPT-4 CoT in GPT-based evaluations, particularly excelling in matchness. These findings underscore ItiNera’s efficacy in enhancing spatial coherence and aligning with user requests in OUIP.
5.4. Ablation Study
To understand how the different components of ItiNera affect the quality of the generated itineraries (RQ2), we compare with the following variants of ItiNera:
  • GPT-4 CoT + UPC: This variant integrates the UPC module to LLMs to generate itineraries conditioned on user-owned POIs.
  • Ours w/o RD: This variant leverages the entire user input string’s embedding to retrieve the POIs.
  • Ours w/o PPR: This variant quantifies the contribution of the PPR module compared to our full system.
  • Ours w/o CSO: This variant removes the CSO module and let the LLM called in the IG module determines the order of candidate POIs when generating the final itineraries.
  • Ours w/ GPT-3.5: This variant replace the GPT-4 with GPT-3.5 for generating the final itinerary.
We remove Fail Rate in ablation study, since all variants equipped with UPC never generate POIs not presented in the database.
The results are shown in Tab. 2. Comparing the first two rows, we find UPC can improve Recall Rate and Matchness of the GPT-4 CoT baseline. The variants “w/o RD”, “w/o PPR”, and “w/ GPT-3.5” exhibit lower Recall Rate, POI Quality and Matchness than our full model. This is because they are relatively inept at understanding user requests and selecting relevant POIs. However, they display lower Average Margin and Overlaps. We attribute this to the idea that there is a trade-off between spatial optimization and matchness. Imposing constraints to align a system with human preference will sacrifice its inherent ability. This is similar to conditional generation tasks in other domains (Saharia et al., 2022; Di et al., 2021) , where aligning output with human input hinders the generation quality and is worth than unconditional generation. This trade-off is also observed when removing CSO module (“w/o CSO”) leads to worse Average Margin and Overlaps but better Recall Rate, POI Quality and Matchness. Our full model manages to balance matchness with spatial ability.
In addition to effectiveness of the PPR module as verified by the results of “w/o PR”, the integration of the PPR module could also help us address the issue of the limited sizes of LLMs’ context windows, saving costs on LLM for our system. Lastly, we find that the results of the variant “w/ GPT3.5”, although compromised, still significantly outperform the GPT-3.5 baseline in Tab. 1, demonstrating our system’s adaptability to alternative LLMs for OUIP.
5.5. Qualitative Results
We further conduct a qualitative study to demonstrate the effectiveness of our system and the importance of integrate LLM with a CSO module. Consider a user request “I’m seeking an artsy itinerary that includes exploring the river’s bridges and ferries.” We visualize the results from ItiNera and GPT-4 CoT in Fig. 3.
Figure 3. Itineraries of ItiNera (left) and GPT-4 CoT (right).
We find that our itinerary better matches the user preferences. The itinerary passes several bridges along the Huangpu River, include a ferry crossing, and conclude at the art-atmosphere-rich Duoyun Bookstore, offering a restful endpoint for users. In contrast, the POIs selected by GPT are more mainstream. Moreover, our spatial arrangement is more logical, avoiding detours and concentrating selected POIs within two spatial clusters. The itinerary generated by GPT is spatially poor, with a disordered sequence of visits and contains excessively distant POIs. Beyond this example, GPT also risks hallucinating non-existent POIs, highlighting the superiority of our system ItiNera in comparison.
5.6. Online Performances
5.6.1. Online Deployment
Our system has been deployed as the core algorithm of the TuTu online OUIP service. Within a month of operation, it has attracted over 1K registered users who have created more than a thousand itineraries across 60 cities, incorporating over 30K user-owned POIs. We provide demonstration of the deployed service in Appendix A.
In our online production, the itinerary generation utilizes the GPT-4 API with streaming. The average latency for the initial response token is around 6 seconds, and the complete response takes around 20 seconds. Our designed prompt asks for a short output, thereby reducing the overall response time.
5.6.2. Online Evaluation
Method Rule-based POI Quality Itinerary Quality Matchness
AM OL Expert User GPT Expert User GPT Expert User GPT
@@ 5.6.2. Online Evaluation
Ours 107.6 0.44 3.8 4.3 70% 3.2 3.8 68% 3.6 3.5 72%
Table 3. Online evaluation.
To verify the effectiveness of our system in real-world scenarios, we conduct online evaluations (RQ3). Subjective evaluation via user studies has been extensively employed in prior research on generative tasks (Saharia et al., 2022; Rombach et al., 2022; Zhuo et al., 2023) where objective metrics fail to adequately assess specific dimensions of the output quality. The TuTu online OUIP service provides a natural way to conduct subjective evaluation. For a small group of users (User), we provide them with a feedback interface upon the completion of each itinerary generation. For those users, we randomly replace the algorithm in the service with the baseline GPT-4 CoT to generate itineraries in the probability of 50%. In the same way as the GPT-based metrics, users are required to rate the POI Quality, Itinerary Quality and Matchness, following the Likert scale (1 for poor and 5 for excellent). We gathered 1443 results from 464 users. These ratings are subsequently utilized to upgrade our system to enhance user experience. We also invited 33 experienced travel assistants (Expert) from our collaborative travel agency to provide evaluation on the collected 1443 queries and itineraries from a professional perspective. For GPT-based metrics, we first use the other algorithm to generate an itinerary for the same query to compare with the existing itinerary. Then, we ask GPT-4 to rank the two candidate itineraries and compute the win rate over all queries.
@@
The average evaluation results are reported in Tab. 3. One can observe that our method is preferred by experts and online users across all metrics, especially for Matchness, validating the effectiveness of our system in real-world scenarios. Moreover, the online evaluation results are consistent with the GPT evaluation win rate, indicating that the proposed GPT-based metrics are appropriate and adaptable where massive results need to be evaluated but rule-based evaluation is not sufficient.
6. Conclusion
@@
We introduce the problem of OUIP and a novel solution approach ItiNera that integrates Large Language Models (LLMs) with spatial optimization. ItiNera enables the generation of personalized and spatially coherent itineraries directly from natural language requests. Our system is deployed in the TuTu online service. The results from offline and online experiments validate the effectiveness of our approach. This study not only establishes a new benchmark for itinerary planning technologies but also broadens venues for further innovations in leveraging LLMs for complex problem-solving in urban contexts. Future studies will focus on categorization of diverse user requests, and also further enhance the system’s scalability and real-time performance, to better serve the growing user base of the TuTu online service.
@@ Acknowledgement
We thank Zhaofeng Wu and Han Zheng at MIT, and Tiange Luo at UMich, for participating in discussions about the initial idea and providing valuable insights on this project.
Appendix A Demonstration of the Deployed Service
Figure 4. Screenshots of the deployed service: POI view & Itinerary view.
We provide screenshots of our deployed system in TuTu online service in Fig. 4. The left screenshot shows the POI interface, where users can add new POIs by direct searching or pasting a link of a travel-related post. They can filter their desired POIs to display on their personal map. The POI icon represent its category (entertainment, restaurant, etc.). By pressing the bottom right button, users can select several POIs to create an itinerary. They can also use our proposed system to generate an itinerary from natural language input (the left figure of Fig. 1).
The right screenshot shows the itinerary interface. Users can browse the itineraries they have created and generated. They can tap one itinerary see the details (the middle and right figures of Fig. 1).
@@ Appendix B Dataset
Table 4. POI data sample.
[…]
@@
Table 5. Itinerary data sample.
The sample itinerary data is shown in Table 5, which contains only two columns: one for the user’s request and the other storing a list of POI IDs that represent the ground truth itinerary (label) for the user’s request.
@@ Appendix C Implementation Details
C.1. Baseline Settings
For the baseline IP, we adopt the $\small\mathbb{P}_{IT}$ for time budgeting. For the LLM baselines, we prompt them to generate itineraries based on user requests. For each POI in the generated itinerary, we search for it using the Map API. Here, the database associated with the Map API is considered to be the current collection of all existing POIs. We leverage fuzzy string matching  [^3] https://github.com/seatgeek/thefuzz to determine if there is a match with specific POIs. The failed POIs contribute to the failure rate metrics. For the matched POIs, attributes of the POI (such as location) are attached to the itinerary for subsequent evaluation.
C.2. Production Environment Details
In our production environment, we leverage the AWS ECS service to orchestrate a cluster of servers, each provisioned with an 8-core CPU and 16GB of RAM, optimizing the performance for Python-based applications and machine learning APIs. To ensure high availability and enhanced data input/output operations, we employ a multi-region Amazon RDS setup, complemented by Redis  [^4] https://redis.io/ for advanced caching strategies. Additionally, we integrate Amazon OpenSearch Service  [^5] https://aws.amazon.com/opensearch-service/ for executing complex queries and supporting sophisticated search and embedding functionalities. To manage user AI API requests efficiently, we’ve implemented a message queue-like system using Amazon Simple Queue Service (SQS), which offers a highly scalable, reliable, and secure means of message queuing and inter-process communication.

### v4

@@ 4.1 Experimental Setup
We collect an urban travel itinerary dataset from four Chinese cities in collaboration with a leading travel agency specializing in single-day citywalk. Each data sample contains a user request, the corresponding urban itinerary plan, and detailed POI data. In total, the dataset covers 1233 top-rated urban itineraries and 7578 POIs. For detailed data format, sample entries, and key preprocessing methodologies employed, refer to §B.
We use GPT-4 for final itinerary generation to ensure quality and GPT-3.5 for other interactions to speed up responses. Our system and data are originally in Chinese, and we provide a translated version in this paper. Additional implementation details are in §C.
Variants UPC RD PPR CSO IG Rule-based Metrics LLM-Eval $\uparrow$ (%)
$R$ $\uparrow$ AM $\downarrow$ OL $\downarrow$ PQ IQ Match
GPT-4 CoT $\times$ $\times$ $\times$ $\times$ 18.4 258.3 0.49 / / /
GPT-4 CoT + UPC $\times$ $\times$ 34.2 240.2 0.52 65.5 61.8 70.6
ItiNera w/o RD $\times$ 22.6 35.4 0.18 68.2 61.5 60.5
ItiNera w/o PPR $\times$ 28.2 84.6 0.38 66.7 63.4 62.2
ItiNera w/o CSO $\times$ 32.8 242.8 1.04 72.1 60.2 74.2
ItiNera w/ GPT-3.5 27.6 79.4 0.56 67.3 58.8 61.4
ItiNera w/ LLaMA 3.1 8B 27.8 90.6 0.45 66.9 58.6 63.5
ItiNera (full) 31.4 86.0 0.42 69.8 64.6 72.0
Table 2: Ablation study on Shanghai dataset.
4.2 Evaluation Metrics
A satisfactory itinerary must be spatially coherent and aligned with the user’s needs, so we designed the following evaluation metrics.
Rule-based Metrics
@@ 4.2 Evaluation Metrics
(1) Recall Rate (RR): the recall rate of POIs in the ground truth itinerary, which evaluates the accuracy of understanding user requests and recommending personalized POIs. (2) Average Margin (AM): the average difference per POI between the total distance of the generated itinerary and the shortest distance (via TSP). (3) Overlaps (OL): the number of self-intersection points in the generated itinerary. AM and OL measure spatial optimization for POI visit order, with lower values being better. (4) Fail Rate (FR): the percentage of POIs from LLM not matched with queried map service POIs, which assesses the LLM’s information accuracy, as failed POIs are inaccessible and impact the user experience.
LLM-Evaluated Metrics
@@
The rule-based metrics are intuitive, but some aspects, like POI appeal and alignment with user requests, are hard to quantify. Thus, we propose several LLM-evaluated metrics: (1) POI Quality (PQ): how interesting and diverse the POIs are; (2) Itinerary Quality (IQ): the overall quality and coherence of the itinerary; (3) Match: the alignment between the itinerary and the user request. We use GPT-4 to rank two itineraries and compute the win rate, repeated at least 10 times for reliability. Our LLM-evaluated metrics have been shown to be consistent with human judgments, as discussed in Sec. 4.5.
4.3 Overall Results
We consider the following baselines:
@@ 4.3 Overall Results
  • IP Gunawan et al. (2014): A traditional IP method. We simplify it to use LLM for time budgeting and to consider POI ratings as utilities.
@@
  • Ernie-Bot 4.0 Sun et al. (2021): The best-performing model on Chinese LLM tasks, selected as our dataset and system are in Chinese.
  • GPT-3.5, GPT-4 and GPT-4 CoT OpenAI (2023): ChatGPT models with or without Chain-of-Thought Wei et al. (2022).
The baseline IP and our method do not compute the Fail Rate since the candidate POIs are all from the dataset.
@@
The result is shown in Tab. 1. our proposed ItiNera outperforms all baselines across all metrics and achieves better or comparable results compared with ground truth data. It shows a $\approx$30% improvement in rule-based metrics over the best baseline, demonstrating superior personalization of user experiences. It maintains spatial coherence, generating itineraries only $\approx$100 meters longer per POI than the shortest TSP-solved path. ItiNera is also the only method to outperform GPT-4 CoT in LLM-evaluated metrics, especially in Match. These results highlight ItiNera’s effectiveness in enhancing spatial coherence and aligning with user requests in OUIP.
Method Rule-based POI Quality Itinerary Quality Match
AM OL Expert User LLM Expert User LLM Expert User LLM
@@
ItiNera 107.6 0.44 3.8 4.3 70% 3.2 3.8 68% 3.6 3.5 72%
Table 3: Deployed System Performance.
4.4 Ablation Study
To validate the effect of each component, we compare the following variants of ItiNera:
  • GPT-4 CoT + UPC: integrates the UPC module to LLMs to generate itineraries based on user-owned POIs.
  • ItiNera w/o RD: uses the entire user input string’s embedding to retrieve POIs.
  • ItiNera w/o PPR: quantifies the contribution of the PPR module compared to our full system.
  • ItiNera w/o CSO: removes the CSO module and lets the LLM in the IG module determine the order of candidate POIs for the final itineraries.
  • ItiNera w/ GPT-3.5 or LLaMA 3.1 8B: replaces GPT-4 with either GPT-3.5 or LLaMA 3.1 8B for generating the final itinerary.
We remove Fail Rate in the ablation study since all variants equipped with UPC never generate POIs not present in the database.
The results in Tab. 2 show that UPC enhances the Recall Rate and Match of the GPT-4 CoT baseline. Variants “w/o RD,” “w/o PPR,” and “w/ GPT-3.5” have lower Recall Rate, POI Quality, Itinerary Quality, and Match than our full model, indicating a trade-off between spatial optimization and alignment with user requests. This parallels other conditional generation tasks, where aligning with human preferences can reduce inherent system ability Saharia et al. (2022); Di et al. (2021). Removing the CSO module worsens the Average Margin and Overlaps but improves Recall Rate, POI Quality, and Match, showing the full model balances alignment with spatial ability. “W/o PPR” shows that the PPR module can address LLM context window limitations and save costs. Finally, “w/ GPT-3.5” outperforms the GPT-3.5 baseline, demonstrating our system’s adaptability to different LLMs.
To validate that our method is a general framework compatible with both open-source LLMs and commercial ones, we conduct experiments with LLaMA 3.1-8B-Instruct Dubey et al. (2024), a state-of-the-art model suitable for consumer GPUs. LLaMA 3.1 offers performance comparable to GPT-3.5. Nevertheless, the performance of open-source models still lags behind commercial models. Considering the maintenance cost of hosting open-source models locally, we opt to use commercial models through API following most companies.
4.5 Deployed System Performance
Our deployed system is currently accessible to a select group of users recommended by our partnered travel agency. To verify the effectiveness of our system in real-world scenarios, we conduct human evaluations. Human evaluation has been extensively employed in prior research on generative tasks Saharia et al. (2022); Rombach et al. (2022); Zhuo et al. (2023) where objective metrics fail to adequately assess specific dimensions of output quality. We invite 464 regular users of our system (User) and 33 experienced travel assistants from our partnered travel agency (Expert) to compare the two itineraries (randomly ordered) generated by GPT-4 CoT and our system based on their requests.
@@ 4.5 Deployed System Performance
The average evaluation results in Tab. 3 show that our method is preferred by both experts and regular users across all metrics, especially for Match, validating the effectiveness of our system in real-world scenarios. The human evaluation results are consistent with the LLM evaluation win rate, indicating that the proposed LLM-evaluated metrics are appropriate and adaptable when rule-based evaluation is insufficient.
4.6 Qualitative Results
We further conduct a qualitative study to demonstrate the importance of integrating LLM with spatial optimization. Consider a user request “I’m seeking an artsy itinerary that includes exploring the river’s bridges and ferries”, we visualize the results from ItiNera and GPT-4 CoT in Fig. 3.
Figure 3: Generated itineraries of ItiNera (left) and GPT-4 CoT (right).
We find that our itinerary better matches the user preferences. The itinerary passes several bridges along the Huangpu River, includes a ferry crossing, and concludes at the art-atmosphere-rich Duoyun Bookstore, offering a restful endpoint for users. In contrast, the POIs selected by GPT are more mainstream. Moreover, our spatial arrangement is more logical, avoiding detours and concentrating selected POIs within two spatial clusters. The itinerary generated by GPT is spatially poor, has a disordered sequence of visits, and contains excessively distant POIs. Beyond this example, GPT also risks hallucinating non-existent POIs, highlighting the superiority of our system in comparison.
5 Conclusion
@@
We introduce the OUIP problem and a solution ItiNera that integrates LLMs with spatial optimization. ItiNera enables the generation of personalized and spatially coherent itineraries directly from natural language requests. Experiments on the real-world dataset and deployed system performance validate the effectiveness of our approach. This study not only sets a new benchmark for itinerary planning technologies but also broadens venues for further innovations in leveraging LLMs for complex problem-solving in urban contexts.
@@ Acknowledgement
We thank Han Zheng at MIT and Tiange Luo at UMich for participating in discussions about the initial idea and providing valuable insights.
Limitations
Despite the success of ItiNera in generating personalized itineraries, our system has several limitations. First, while the spatial optimization module works well in many cases, it may face efficiency challenges in highly complex urban environments. Moreover, although LLMs provide significant language processing capabilities, they still exhibit limitations in spatial reasoning and real-time decision-making, which may impact the quality of the generated itineraries in specific scenarios.
Ethical Statement
This study adheres to strict ethical guidelines to protect user privacy and data. The personalized POI database is user-owned, and all data processing follows legal data security and user consent standards. We have designed the system to be fair and inclusive, avoiding biases in itinerary recommendations across diverse user groups. Additionally, we emphasize environmental responsibility by ensuring that the system promotes sustainable urban tourism without adversely affecting local culture or ecosystems. Finally, transparency in the pipeline is a priority, ensuring users understand how their itineraries are generated.
Appendix A Demonstration of the Deployed System
Figure 4: Screenshots of the deployed system: POI view & Itinerary view.
We provide screenshots of our deployed system in Fig. 4. The left screenshot shows the POI interface, where users can add new POIs by direct searching or pasting a link of a travel-related post. They can filter their desired POIs to display on their personal map. The POI icon represents its category (entertainment, restaurant, etc.). Users can select several POIs by pressing the bottom right button to create an itinerary. They can also use our system to generate an itinerary from natural language input (the left figure of Fig. 1).
The right screenshot shows the itinerary interface. Users can browse the itineraries they have created and generated. They can tap one itinerary to see the details (the middle and right figures of Fig. 1).
@@ Appendix B Dataset
Table 4: POI data sample.
The sample POI data is shown in Table 4, where the context column is a concatenation of the strings from all the previous columns. The embedding of each POI is also obtained by calling $E_{\theta^{\prime}}$ to embed the context field. The resulting embedding, $\mathcal{E}$, contains rich semantic information about the POIs.
@@
Table 5: Itinerary data sample.
The sample itinerary data is shown in Table 5, which contains two columns: one for the user’s request and the other storing a list of POI IDs representing the ground truth itinerary (label) for the user’s request.
@@ Appendix C Implementation Details
C.1 Method Implementation
We use the OpenAI text-embedding-ada-002 model for embedding purposes. The spatial coherence of itineraries is optimized through an open-source TSP solver [^1] https://github.com/fillipe-gsm/python-tsp. Integration of POI data, including geographical coordinates, user ratings, categorizations, and physical addresses, is facilitated through the Amap API [^2] https://lbs.amap.com/.
C.2 Baseline Settings
We use the same itinerary generation prompt for all baselines, including basic task requirements and output format, as in $\mathbb{P}_{IG}$ in §F.4. For GPT-4 CoT, we extend the prompt by integrating “thoughts”, detailed in §F.5.
We adopt the $\mathbb{P}_{IT}$ for the baseline IP for time budgeting. We prompt the LLM baselines to generate itineraries based on user requests. We searched for each POI in the generated itinerary using the Map API. Here, the database associated with the Map API is considered to be the current collection of all existing POIs. We leverage fuzzy string matching [^3] https://github.com/seatgeek/thefuzz to determine if there is a match with specific POIs. The failed POIs contribute to the failure rate metrics. For the matched POIs, attributes of the POI (such as location) are attached to the itinerary for subsequent evaluation.